OpenStudy (anonymous):

square root of (2x) times cube root of (3x^2)

OpenStudy (anonymous):

root(2)*cuberoot(3)*x^2

OpenStudy (anonymous):

\[\sqrt{2x}\times \sqrt[3]{3^2}\]